Simplify a trigonometric expression sin θ cos2 θ − sin θ


1
Expert's answer
2021-11-30T16:10:15-0500

"sin\theta cos2\theta - sin\theta \\

\text{Recall that} \, cos2\theta = 1 - 2sin^2\theta. \text{Hence, we have}\\

sin\theta(1 - 2sin^2\theta) - sin\theta \\
= sin\theta - 2sin^3\theta - sin\theta\\
= -2sin^3\theta"